# Learning in Compact Spaces with Approximately Normalized Transformers

### v1

@@ Acknowledgements
We gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this work by providing computing time through the John von Neumann Institute for Computing (NIC) on the supercomputer JUWELS Booster at Jülich Supercomputing Centre (JSC), EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LEONARDO, hosted by CINECA (Italy) and the LEONARDO consortium through an EuroHPC Extreme Access grant EHPC-EXT-2023E02-068, storage resources on JUST granted and operated by JSC and supported by Helmholtz Data Federation (HDF), compute resources provided via WestAI compute grant "Measuring and enhancing advanced reasoning capabilities of foundation models via local model deployment" (westai0007) at JSC and compute resources provided via JSC at JURECA.
@@ Appendix A Related Work
Previous approaches have addressed the “Curse of Depth” [6] through depth-specific scaling [25], initialization strategies [26], or hybrid normalization schemes [27]. Sun et al. [6] introduced a constant scaling factor depending on the layer number after the pre-normalization layer. This reduces the effect of increasing variance on the residual, but in contrast to our approach, it does not aim to eliminate the effect. In this work, we propose a comprehensive solution that normalizes the entire representation at each layer, ensuring all network components contribute effectively regardless of depth.
Franke et al. [13] introduce Constrained Parameter Regularization which bounds a statistical measure, like the $L_{2}$-norm, of learnable parameters using an augmented Lagrangian optimization instead of applying weight decay. Therefore, they introduce multiple initialization methods to find the right norm value. In contrast, our work bound all parameter matrices to one, and we also approximately normalized the representation space.
Multiple works proposed methods or architecture changes to remove or replace the normalization [28, 29, 30, 31]. Most recently, Zhu et al. [32] proposed to replace the normalization layer with a Dynamic Tanh layer (DyT) based on the observation that normalization produces S-shaped input-output mappings. DyT consists of a tanh function with an input scaling scalar and output scaling vector. In contrast to our approach, DyT does not change the representation or weight space and only aims for an improved runtime.
@@
We compare our anGPT approach to the constant scaling factor [6] and the dynamic tanh (DyT) replacement of the layer normalization [32]. We train both on a $0.5B$ GPT setting on $10B$ SlimPajama tokens (Chinchilla optimal) using the same configuration as for GPT+. However, we tune the learning rate for both approaches. We apply the additional learnable scaling factor for DyT as proposed and a $\alpha_{0}$ initialization of $1.0$ according to Table 12 in [32]. We include in the comparison nGPT and GPT+ and show the results in Figure A.1.
@@
Figure A.1: Comparison of GPT+, nGPT, and anGPT to the LN scaling [6] and DyT normalization [32] in a 0.5B GPT training on 10B SlimPajama token budget. We use the same configuration as for GPT+ and tune the learning rate.
@@ Appendix B Variance on the Adam momentum
Appendix C RMSNorm scaling norm and block-input norm in GPT+
@@ Appendix D Normalizing the Attention Matrix
To apply the normalization for the linear map to $\bm{A}\bm{v}$, we require the rows $\bm{a}_{i}$ of $\bm{A}$ to be normalized. Assuming the entries $a_{ij}$ of $\bm{A}\in\mathbb{R}^{s\times s}$ are identically distributed, and the softmax computation is applied along the rows $\bm{A}$, we know that $\sum_{j}a_{ij}=1$. By symmetry (identical distribution), we know that $\mathbb{E}[a_{ij}]=1/s$. For the rows $\bm{a}_{i}$ to be normalized, we require $\|\bm{a}_{i}\|_{2}^{2}=\sum_{j}a_{ij}^{2}=1$, i.e., $\mathbb{E}[a_{ij}]=1/\sqrt{s}$. We therefore scale the attention matrix by $s/\sqrt{s}=\sqrt{s}$. If a mask for causal attention is used, we know the number of zero entries in each row. Thus, the expected value of the nonzero entries in the $r$-th row is $\mathbb{E}[a_{rj}]=1/r$, so the $r$-th row is scaled by $r/\sqrt{r}=\sqrt{r}$.
Note that the normalization factors are likely less useful for attention as the expected values effectively model dense attention scores, while sparse attention is usually more realistic. Hence, calculating useful normalization factors for the attention matrix likely requires information about the number of effectively nonzero attention scores. Unfortunately, this information is often not directly accessible due to the use of FlashAttention [33]. Approaches to leverage this information for effective normalization may be seen as directions for future work.
Appendix E Norms of the parameter groups in anGPT
Figure E.1: The mean norm of dimension 1 (input dimension) of anGPT parameter groups during training. The traces correspond to individual layers. We see that despite the bound, the norm stays close to one.
@@ Appendix F Training Details
In our experiments, we use model sizes from $32M$ up to $1B$ parameters and list the architecture hyperparameters in Table F.1. For the optimization, we use Adam and AdamW and list the corresponding training hyperparameters in Table F.2. We used two datasets in this paper, SlimPajama (Apache 2.0 license) [14] and OpenWebText (Creative Commons Zero v1.0) [24] [^1] Both datasets are accessible on Huggingface: https://huggingface.co/datasets/cerebras/SlimPajama-627B and https://huggingface.co/datasets/Skylion007/openwebtext. SlimPajama provides a validation set, and for OpenWebText, we used 10k randomly selected documents as a validation set.
We implemented our experiments in PyTorch 2.6 [34] and used Flash Attention 0.7.3 [33]. All plots are generated with Matplotlib [35]. We performed all experiments on a research cluster with 4×4$\times$4 × A100 40GB GPU nodes and used in total about 30k GPU hours. The smallest experiments are around 1 GPU hour, and the largest are up to 750 GPU hours.
@@
Number of Attn. Heads ($n_{h}$) 4 6 8 12 16 20
Head Dim. ($d_{k}=d_{m}/n_{h}$) 64 64 64 64 64 64
MLP Dim. ($d_{MLP}=4\times d_{m}$) 1024 1536 2048 3072 4096 5120
@@
Parameter GPT+ nGPT anGPT
@@
Optimizer AdamW Adam Adam
@@
Weigth decay 0.1 0 0
Lr Num warm-up Steps 20% 0 0
@@
Param. Scale Init - $1/\sqrt{d_{m}}$ 0.01
@@ Appendix G Fitting hyperparameter scaling trends
To investigate the optimal hyperparameters and scaling trends of the new anGPT architecture, we performed a grid search on different scales and used the extrapolated optimal configuration to train multiple models on different token budgets. We orient our procedure on Porian et al. [21] work investigating the discrepancies in compute-optimal scaling of language models between Kaplan et al. [36] and Hoffmann et al. [22].
For the grid search, we performed training runs with GPT+ and anGPT on the SlimPajama dataset with different model scales from $32M$ to $0.5B$ parameters. We train a Chinchilla optimal token budget of 20×20$\times$20 × the number of training parameters [36]. Similar to Porian et al. [21], we used an Adam $\beta_{2}$ parameter of $0.99$ for experiments below $100M$ parameters and $0.95$ above. We performed at least three experiments per scale and batch size with different learning rates, so that the best configuration is always in the middle of the parameter grid. Our raw results of the hyperparameter sweeps can be found in Figure G.1.
@@
Our interpolation method for estimating the optimal batch size and learning rate closely follows the two-stage procedure proposed by Porian et al. [21]. In their method in the first stage, for each model size and batch size, the optimal learning rate was identified by performing Akima interpolation (in log-space) on the loss as a function of the learning rate, taking the lowest loss among three tested values of the hyperparameter $\beta_{2}$, and subsequently identifying the minimizing argument. In the second stage, they applied interpolation again, this time over batch sizes, using the previously interpolated minimal losses to pinpoint an optimal batch size. The final optimal learning rate for the identified batch size was obtained by interpolating the sequence of (batch size, minimizing learning rate) pairs and evaluating this interpolant at the determined optimal batch size.
@@
Figure G.1: Hyperparameter sweep results: The validation loss after $20\times N$ training steps as a function of the learning rate for different model sizes $N$ and batch sizes. Plot design inspired by Wortsman et al. [37]. The stars with a black outline indicate the interpolated minimum learning rate as used in the first stage of the estimation of the optimal batch size and learning rate.
@@ Appendix H Comparison to nGPT
Since we compare our work with nGPT [5], we performed a sanity check with our code base and compared the results of the official nGPT implementation [38]. We reconstructed the values of the reported model performance across different learning rates from the repository in Figure H.1 and added our results. We find the final validation loss values for nGPT and GPT without QK norm match the reported numbers. We also found anGPT has a slightly lower final validation loss ($0.6\%$) than nGPT, but, maybe more importantly, we also found that GPT with QK norm increases the GPT baseline performance substantially ($2.6\%$ lower final validation loss).
@@
This finding is in line with our experiment on SlimPajama, shown in Figure H.2. We find a higher convergence speedup when compared to GPT+ without QK norm. However, there is still a gap to the reported speedup factors of 4×4$\times$4 × up to 20×20$\times$20 × by Loshchilov et al. [5]. We hypothesize this could be due to different training data, training budget, hyperparameters, and/or the codebase.
In contrast to nGPT, we use a larger training corpus to perform LLM pretraining experiments without training multiple epochs. While Loshchilov et al. [5] used OpenWebText [24] with ~$\text{\textasciitilde}9B$ tokens (32k tokenizer) and trained for multiple epochs (up to 50 epochs), we use SlimPajama [14] with ~$\text{\textasciitilde}627B$ tokens (50k tokenizer) and train for $<1$ epoch. Also, we trained only up to ×$\times 7$ Chinchilla optimal [22] token budgets while nGPT was trained on up to ×$\times 20$ Chinchilla optimal token budgets. Furthermore, we tuned the batch size and found a smaller batch size for GPT+ slightly better than for anGPT. We trained only on a context length of 2k tokens (since the average document length in SlimPajama is only  1k tokens) while nGPT was trained on up to 8k tokens context size. Lastly, the publicly available codebase on GitHub is different from the one used in the paper, as the author explains in GitHub Issue 6 [38]. Nevertheless, the nGPT codebase was very helpful for our work.
Figure H.2: Training the 0.5B model up to 7×7$\times$7 × Chinchilla optimal token budget. Each point is a full training with the training budget noted on the abscissa. The plot below shows the convergence speed up against the GPT+ model without QK normalization.
Appendix I Derive scaling laws
For each combination of compute scale $C$ and architecture, we select a point with the minimal validation loss. We use the approach described in Hoffmann et al. [22]: we bin compute into 1500 FLOPs logarithmically spaced intervals and for each bin we obtain a point with the minimal loss. We obtain a mapping from each combination of parameters and number of tokens to the compute C. Following Hoffmann et al. [22], [36] we assume that the validation loss $\mathcal{L}(C)$ is proportional to $C^{-\alpha}$ ($\alpha>0)$ up to some positive constant $A_{0}$.
@@ Appendix I Derive scaling laws
We follow [22] and model compute optimal number of tokens $N_{opt}(C)$ and $D_{opt}(C)$ as power laws. From the set of point $\{N,D\}$ we select such $D$ and $N$ that correspond to minimal validation loss:
@@
where $t_{\alpha/2,\,n-p}$ is the critical value from the Student’s $t$-distribution at a significance level of $\alpha=0.05$. We fit power functions through the obtained points to estimate parameters for $D_{opt}(C)$ and $N_{opt}(C)$. The obtained power law fits are shown in Figure I.1, and the estimated parameters are in Table I.1.
@@
Table I.1: Estimated scaling law parameters for compute-optimal dataset size for $D_{\text{opt}}(C)$ and compute-optimal model size $N_{\text{opt}}(C)$. The table reports fitted values for both anGPT and GPT+. We observe that both models exhibit very similar exponents, which suggests comparable trends in how optimal dataset and model size should be selected under certain compute.
To determine whether estimated parameters (exponents $\alpha_{GPT+}$ and $\alpha_{anGPT}$ for GPT+ and anGPT fits respectively) are statistically different, we construct a Z-test:
  • $H0$: $\alpha_{GPT+}=\alpha_{anGPT}$.
  • $H1$: $\alpha_{GPT+}\neq\alpha_{anGPT}$.
Using covariance matrix of parameters $\text{Cov}(\hat{\theta})$ we obtain variances $\sigma_{GPT+}^{2}$ for $\alpha_{GPT+}$ and $\sigma_{anGPT}^{2}$ for $\alpha_{anGPT}$. Then we can use these variances to obtain a z-score and the two sided p-value. If the obtained $\textit{p}<\alpha$ ($\alpha=0.05$ for $95$% confidence interval) we must reject the null hypothesis. We obtain $\sigma_{anGPT}^{2}=1.11\cdot 10^{-6}$ and $\sigma_{GPT+}^{2}=1.08\cdot 10^{-6}$. Therefore, we get $p=0.073>\alpha$ which means that we cannot reject the null hypothesis and the estimated exponents $\alpha_{anGPT}$ and $\alpha_{anGPT}$ are not significantly different.
(a) Compute optimal number parameters $N_{opt}(C)$
(b) Compute optimal dataset Size $D_{opt}(C)$
@@
Figure I.1: Scaling laws for compute-optimal model size and dataset size. a) shows the estimation for compute optimal number of parameters $N_{opt}(C)$ and b) shows compute-optimal dataset size $D_{opt}(C)$, both as function of compute $C$. We observe that both anGPT and GPT+ have similar scaling trends for compute-optimal allocations. Notably, both fits exhibit high uncertainty, which is indicated by the wide confidence intervals surrounding the curves.

### v2

@@ Acknowledgements
We thank Timo Denk, Aaron Klein and Frederic Runge for valuable discussions and feedback on the manuscript. Their insights greatly improved the clarity and quality of this work.
@@
We gratefully acknowledge the Gauss Centre for Supercomputing e.V. for funding this work by providing computing time through the John von Neumann Institute for Computing (NIC) on the supercomputer JUWELS Booster at Jülich Supercomputing Centre (JSC), EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LEONARDO, hosted by CINECA (Italy) and the LEONARDO consortium through an EuroHPC Extreme Access grant EHPC-EXT-2023E02-068, storage resources on JUST granted and operated by JSC and supported by Helmholtz Data Federation (HDF), compute resources provided via WestAI compute grant "Measuring and enhancing advanced reasoning capabilities of foundation models via local model deployment" (westai0007, westai0066) at JSC and compute resources provided via JSC at JURECA.
@@ Appendix A Related Work
Previous approaches have addressed the “Curse of Depth” [sun2025curse] through depth-specific scaling [bachlechner21a], initialization strategies [zhang2018residual], or hybrid normalization schemes [liu2020understanding]. sun2025curse introduced a constant scaling factor depending on the layer number after the pre-normalization layer. This reduces the effect of increasing variance on the residual, but in contrast to our approach, it does not aim to eliminate the effect. In this work, we propose a comprehensive solution that normalizes the entire representation at each layer, ensuring all network components contribute effectively regardless of depth.
franke2024improving introduces Constrained Parameter Regularization which bounds a statistical measure, like the $L_{2}$-norm, of learnable parameters using an augmented Lagrangian optimization instead of applying weight decay. Therefore, they introduce multiple initialization methods to find the right norm value. In contrast, our work bound all parameter matrices to one, and we also approximately normalized the representation space.
Multiple works proposed methods or architecture changes to remove or replace the normalization [brock2021characterizing, brock2021high, he2024simplifying, heimersheim2025you]. Most recently, zhu2025transformers proposed to replace the normalization layer with a Dynamic Tanh layer (DyT) based on the observation that normalization produces S-shaped input-output mappings. DyT consists of a tanh function with an input scaling scalar and output scaling vector. In contrast to our approach, DyT does not change the representation or weight space and only aims for an improved runtime.
@@
We compare our anGPT approach to the constant scaling factor [sun2025curse] and the dynamic tanh (DyT) replacement of the layer normalization [zhu2025transformers]. We train both on a $0.5B$ GPT setting on $10B$ SlimPajama tokens (Chinchilla optimal) using the same configuration as for GPT+. However, we tune the learning rate for both approaches. We apply the additional learnable scaling factor for DyT as proposed and a $\alpha_{0}$ initialization of $1.0$ according to Table 12 in [zhu2025transformers]. We include in the comparison nGPT and GPT+ and show the results in Figure A.1.
@@
Figure A.1: Comparison of GPT+, nGPT, and anGPT to the LN scaling [sun2025curse] and DyT normalization [zhu2025transformers] in a 0.5B GPT training on 10B SlimPajama token budget. We use the same configuration as for GPT+ and tune the learning rate.
@@ Appendix B Variance on the Adam momentum
Appendix C RMSNorm $\gamma$-norms and block-input norms in GPT+
@@ Appendix D Normalizing the Attention Matrix
To apply the normalization for the linear map to $\bm{A}\bm{v}$, we require the rows $\bm{a}_{i}$ of $\bm{A}$ to be normalized. Assuming the entries $a_{ij}$ of $\bm{A}\in\mathbb{R}^{s\times s}$ are identically distributed, and the softmax computation is applied along the rows $\bm{A}$, we know that $\sum_{j}a_{ij}=1$. By symmetry (identical distribution), we know that $\mathbb{E}[a_{ij}]=1/s$. For the rows $\bm{a}_{i}$ to be normalized, it is required that $\|\bm{a}_{i}\|_{2}^{2}=\sum_{j}a_{ij}^{2}=1$, i.e., $\mathbb{E}[a_{ij}]=1/\sqrt{s}$. We therefore scale the attention matrix by $s/\sqrt{s}=\sqrt{s}$. If a mask for causal attention is used, we know the number of zero entries in each row. Thus, the expected value of the nonzero entries in the $r$-th row is $\mathbb{E}[a_{rj}]=1/r$, so the $r$-th row is scaled by $r/\sqrt{r}=\sqrt{r}$.
Note that the normalization factors are likely less useful for attention as the expected values effectively model dense attention scores, while sparse attention is usually more realistic. Hence, calculating useful normalization factors for the attention matrix likely requires information about the number of effectively nonzero attention scores. Unfortunately, this information is often not directly accessible due to the use of FlashAttention [dao2022flashattention]. Approaches to leverage this information for effective normalization may be seen as directions for future work.
Appendix E Explicit Normalization Factors
We derive the normalization factors using $\nu_{g}=1/\sqrt{\mathbb{E}[\|g(x)\|_{2}^{2}]}$ for linear maps and Monte Carlo estimation for activation functions. The specific factors are:
  • Query, Key, Value projections: $\nu_{qkv}=\sqrt{d/h}$ for $W_{qkv}\in\mathbb{R}^{h\times d}$
  • Output projection: $\nu_{p}=1$ for $W_{p}\in\mathbb{R}^{d\times d}$
  • Up projection: $\nu_{uz}=\sqrt{d/f}$ for $W_{uz}\in\mathbb{R}^{f\times d}$
  • Activation function: $\nu_{acf}=3.74$ (via Monte Carlo with $10^{5}$ samples)
  • Down projection: $\nu_{d}=\sqrt{f/d}$ for $W_{d}\in\mathbb{R}^{d\times f}$
  • Residual interpolation: $\nu(\alpha)=1/\sqrt{\alpha^{2}+(1-\alpha)^{2}}$
where $d$ is the model dimension, $h$ is the head dimension (64 in our experiments), and $f$ is the feed-forward dimension (typically $4d$).
@@ Appendix F Training Details
In our experiments, we use model sizes from $32M$ up to $1B$ parameters and list the architecture hyperparameters in Table F.1. For the optimization, we use Adam and AdamW and list the corresponding training hyperparameters in Table F.2. We used two datasets in this paper, SlimPajama (Apache 2.0 license) [cerebras2023slimpajama] and OpenWebText (Creative Commons Zero v1.0) [Gokaslan2019OpenWeb] [^1] Both datasets are accessible on Huggingface: https://huggingface.co/datasets/cerebras/SlimPajama-627B and https://huggingface.co/datasets/Skylion007/openwebtext. SlimPajama provides a validation set, and for OpenWebText, we used 10k randomly selected documents as a validation set.
We implemented our experiments in PyTorch 2.6 [pytorch-neurips19a] and used Flash Attention 0.7.3 [dao2022flashattention]. All plots are generated with Matplotlib [matplotlib]. We performed all experiments on a research cluster with $4\times$ A100 40GB GPU nodes and used in total about 30k GPU hours. The smallest experiments are around 1 GPU hour, and the largest are up to 750 GPU hours. An open-source implementation of anGPT is available at https://github.com/automl/anGPT.
@@
Number of Attn. Heads ($n_{heads}$) 4 6 8 12 16 20
Head Dim. ($d_{k}=d_{model}/n_{heads}$) 64 64 64 64 64 64
MLP Dim. ($d_{MLP}=4\times d_{model}$) 1024 1536 2048 3072 4096 5120
@@
Parameter GPT+ nGPT/anGPT
@@
Optimizer AdamW Adam
@@
Weigth decay 0.1 0
Lr Num warm-up Steps 20% 0
@@
Param. Scale Init - $1/\sqrt{d_{m}}$ / 0.001
@@ Appendix G Fitting hyperparameter scaling trends
To investigate the optimal hyperparameters and scaling trends of the new anGPT architecture, we performed a grid search on different scales and used the extrapolated optimal configuration to train multiple models on different token budgets. We orient our procedure on porian2024resolving work investigating the discrepancies in compute-optimal scaling of language models between kaplan2020scaling and hoffmann2022compute.
For the grid search, we performed training runs with GPT+ and anGPT on the SlimPajama dataset with different model scales from $32M$ to $0.5B$ parameters. We train a Chinchilla optimal token budget of $20\times$ the number of training parameters [kaplan2020scaling]. Similar to porian2024resolving, we used an Adam $\beta_{2}$ parameter of $0.99$ for experiments below $100M$ parameters and $0.95$ above. We performed at least three experiments per scale and batch size with different learning rates, so that the best configuration is always in the middle of the parameter grid. Our raw results of the hyperparameter sweeps can be found in Figure G.1.
@@
Our interpolation method for estimating the optimal batch size and learning rate closely follows the two-stage procedure proposed by porian2024resolving. In their method in the first stage, for each model size and batch size, the optimal learning rate was identified by performing Akima interpolation (in log-space) on the loss as a function of the learning rate, taking the lowest loss among three tested values of the hyperparameter $\beta_{2}$, and subsequently identifying the minimizing argument. In the second stage, they applied interpolation again, this time over batch sizes, using the previously interpolated minimal losses to pinpoint an optimal batch size. The final optimal learning rate for the identified batch size was obtained by interpolating the sequence of (batch size, minimizing learning rate) pairs and evaluating this interpolant at the determined optimal batch size.
@@
Figure G.1: Hyperparameter sweep results: The validation loss after $20\times N$ training steps as a function of the learning rate for different model sizes $N$ and batch sizes. Plot design inspired by wortsman2024smallscale. The stars with a black outline indicate the interpolated minimum learning rate as used in the first stage of the estimation of the optimal batch size and learning rate.
@@ Appendix H Comparison to nGPT
Since we compare our work with nGPT [loshchilov2025ngpt], we performed a sanity check with our code base and compared the results of the official nGPT implementation [loshchilov2025ngptgithub]. We reconstructed the values of the reported model performance across different learning rates from the repository in Figure H.1 and added our results. We find the final validation loss values for nGPT and GPT without QK norm match the reported numbers. We also found anGPT has a slightly lower final validation loss ($0.6\%$) than nGPT, but, maybe more importantly, we also found that GPT with QK norm increases the GPT baseline performance substantially ($2.6\%$ lower final validation loss).
@@
This finding is in line with our experiment on SlimPajama, shown in Figure H.2. We find a higher convergence speedup when compared to GPT+ without QK norm. However, there is still a gap to the reported speedup factors of $4\times$ up to $20\times$ by loshchilov2025ngpt. We hypothesize this could be due to different training data, training budget, hyperparameters, and/or the codebase.
In contrast to nGPT, we use a larger training corpus to perform LLM pretraining experiments without training multiple epochs. While loshchilov2025ngpt used OpenWebText [Gokaslan2019OpenWeb] with $\text{\textasciitilde}9B$ tokens (32k tokenizer) and trained for multiple epochs (up to 50 epochs), we use SlimPajama [cerebras2023slimpajama] with $\text{\textasciitilde}627B$ tokens (50k tokenizer) and train for <1<1 epoch. Also, we trained only up to $\times 7$ Chinchilla optimal [hoffmann2022compute] token budgets while nGPT was trained on up to $\times 20$ Chinchilla optimal token budgets. Furthermore, we tuned the batch size and found a smaller batch size for GPT+ slightly better than for anGPT. We trained only on a context length of 2k tokens (since the average document length in SlimPajama is only  1k tokens) while nGPT was trained on up to 8k tokens context size. Lastly, the publicly available codebase on GitHub is different from the one used in the paper, as the author explains in GitHub Issue 6 [loshchilov2025ngptgithub]. Nevertheless, the nGPT codebase was very helpful for our work.
Figure H.2: Training the 0.5B model up to $7\times$ Chinchilla optimal token budget. Each point is a full training with the training budget noted on the abscissa. The plot below shows the convergence speed up against the GPT+ model without QK normalization.
Appendix I Convergence Analysis Across Model Scales
Figure I.1 shows convergence comparisons for 250M and 1B parameter models, demonstrating that the  40% speedup observed for 0.5B models (Figure 3 in main paper) is consistent across scales.
Figure I.1: Convergence comparison for (a) 250M and (b) 1B parameter models. Similar to the 0.5B results, anGPT achieves the same loss with significantly fewer iterations, demonstrating consistent  40% speedup across model scales.
Appendix J Derive scaling laws
For each combination of compute scale $C$ and architecture, we select a point with the minimal validation loss. We use the approach described in hoffmann2022compute: we bin compute into 1500 FLOPs logarithmically spaced intervals and for each bin we obtain a point with the minimal loss. We obtain a mapping from each combination of parameters and number of tokens to the compute C. Following hoffmann2022compute, [kaplan2020scaling] we assume that the validation loss $\mathcal{L}(C)$ is proportional to $C^{-\alpha}$ ($\alpha>0)$ up to some positive constant $A_{0}$.
@@ Appendix I Convergence Analysis Across Model Scales
We follow [hoffmann2022compute] and model compute optimal number of tokens $N_{opt}(C)$ and $D_{opt}(C)$ as power laws. From the set of point $\{N,D\}$ we select such $D$ and $N$ that correspond to minimal validation loss:
@@
where $t_{\alpha/2,\,n-p}$ is the critical value from the Student’s $t$-distribution at a significance level of $\alpha=0.05$. We fit power functions through the obtained points to estimate parameters for $D_{opt}(C)$ and $N_{opt}(C)$. The obtained power law fits are shown in Figure J.1, and the estimated parameters are in Table J.1.
@@
Table J.1: Estimated scaling law parameters for compute-optimal dataset size for $D_{\text{opt}}(C)$ and compute-optimal model size $N_{\text{opt}}(C)$. The table reports fitted values for both anGPT and GPT+. We observe that both models exhibit very similar exponents, which suggests comparable trends in how optimal dataset and model size should be selected under certain compute.
((a))
((b))
@@
Figure J.1: Scaling laws for compute-optimal model size and dataset size. a) shows the estimation for compute optimal number of parameters $N_{opt}(C)$ and b) shows compute-optimal dataset size $D_{opt}(C)$, both as function of compute $C$. We observe that both anGPT and GPT+ have similar scaling trends for compute-optimal allocations. Notably, both fits exhibit high uncertainty, which is indicated by the wide confidence intervals surrounding the curves.
Appendix K Downstream Task Evaluation
We evaluate our models on standard benchmarks to verify that pretraining improvements transfer to downstream tasks. Table K.1 shows results for 1B models across different training budgets.
Table K.1: Downstream evaluation on 1B models. Higher is better for all metrics except perplexity (PPL).
Metric Model Training Budget
10B Tokens 21B Tokens 42B Tokens
PIQA (Acc) GPT+ 0.669 0.674 0.700
anGPT 0.680 0.702 0.718
ARC-Easy (Acc) GPT+ 0.508 0.532 0.561
anGPT 0.539 0.561 0.596
HellaSwag (Acc) GPT+ 0.338 0.363 0.400
anGPT 0.354 0.386 0.413
LAMBADA (PPL) GPT+ 36.915 24.932 16.726
anGPT 28.780 17.577 14.679
WikiText (PPL) GPT+ 22.648 21.661 17.621
anGPT 20.518 17.734 16.175
WinoGrande (Acc) GPT+ 0.522 0.536 0.554
anGPT 0.530 0.559 0.563
MMLU (Acc) GPT+ 0.232 0.239 0.243
anGPT 0.232 0.241 0.256
anGPT consistently outperforms GPT+ across benchmarks, with particularly strong improvements in perplexity-based metrics (LAMBADA, WikiText) and reasoning tasks (ARC-Easy, HellaSwag). The performance gains correlate well with the validation loss improvements observed during pretraining.
Appendix L Ablation Studies
We conduct comprehensive ablation studies to investigate the sensitivity of anGPT to variations in normalization factors. Table L.1 shows the impact of different modifications to the normalization scheme on a 0.5B model trained on 10B tokens.
Table L.1: Ablation study on normalization factors. We report relative change to anGPT in parentheses (%).
Configuration Validation Loss PPL
GPT+ 2.677 (+1.46%) 14.541 (+3.94%)
anGPT (baseline) 2.638 13.990
anGPT (const. norm factor $\times$0.5) 2.647 (+0.34%) 14.116 (+0.90%)
anGPT (const. norm factor $\times$2.0) 2.642 (+0.12%) 14.033 (+0.31%)
anGPT (all norm factors $\times$0.5) 7.935 (+200%) 2792.69 (+19k%)
anGPT (all norm factors $\times$2.0) NaN NaN
anGPT (no const. norm factors) 2.667 (+1.09%) 14.400 (+2.93%)
anGPT (no residual norm factor) 2.719 (+3.04%) 15.156 (+8.33%)
anGPT (no norm factors) 2.718 (+3.02%) 15.148 (+8.28%)
anGPT (no LERP) 2.688 (+1.89%) 14.700 (+5.08%)
anGPT (half token budget) 2.755 (+4.42%) 15.718 (+12.36%)
anGPT (double token budget) 2.565 (-2.78%) 13.000 (-7.07%)
These results demonstrate that while the normalization factors are sensitive to large perturbations (50% or 200% scaling), the method remains robust to smaller estimation errors. The residual normalization factors are most critical, as their manipulation can cause training collapse. The 1.09% performance drop from removing constant normalization factors is significant when compared to the performance gap between anGPT and GPT+ (1.46%).
Appendix M Evolution of Learnable Parameters
We track the evolution of learnable interpolation parameters $\alpha_{A}$ (attention) and $\alpha_{M}$ (MLP) throughout training. Figure M.1 shows these parameters across different layers and model sizes. The increase in $\alpha$ values demonstrates that the model learns to increasingly incorporate new features from each block.
Figure M.1: Evolution of interpolation parameters during training of a 0.5B parameter model.
Appendix N Empirical Validation of Norm Concentration
To verify that approximate normalization maintains stable norms throughout the network, we measure the ratio of mean input norm to mean output norm for each block. Table N.1 shows these ratios at different layers.
Table N.1: Ratio of mean output norm to mean input norm for transformer blocks. Values close to 1.0 indicate stable norm propagation.
Layer anGPT GPT+
First attention 1.07 0.01
First MLP 1.86 1.60
Middle attention 0.86 2.47
Middle MLP 0.97 2.60
Last attention 0.92 4.62
Last MLP 1.22 5.83
[…]
Table N.2: Estimation error of normalization factors
Factor Relative Error
$\nu_{qkv}$ 1.16%
$\nu_{p}$ 0.07%
$\nu_{uz}$ 0.02%
$\nu_{d}$ 0.08%
$\nu_{acf}$ 0.41%
$\nu(\alpha)$ 0.00%
All errors are below 2%, well within the robustness margins demonstrated in our sensitivity analysis.
Appendix O Robustness to Distribution Shift
We evaluate norm stability under out-of-distribution inputs by comparing normal text samples against pathological inputs (512 repetitions of the same token). This tests whether fixed normalization factors remain effective under extreme distribution shifts.
Table O.1: Mean residual norms under distribution shift (0.5B model)
Text Input Same Token
Model Mean Std Mean Std
GPT+ 1831.42 353.93 545.68 78.80
anGPT 2.48 0.18 1.65 0.15
[…]
Appendix P Norms of the parameter groups in anGPT
Figure P.1: The mean norm of dimension 1 (input dimension) of anGPT parameter groups during training. The traces correspond to individual layers. We see that despite the bound, the norm stays close to one.